OpenStudy (anonymous):

What is the x-coordinate of the x-intercept of the line that passes through points (-5, 3) and (1, -3)?

OpenStudy (anonymous):

You want to find the point (x,0) (the point which intercepts the x axis). To do so, you must find the linear equation of the function which is going through these points, and then you plug y=0. Can you manage ?

OpenStudy (anonymous):

So how would I find the linear equation of the function then given the given information?

OpenStudy (anonymous):

A linear equation is of this form: y=ax+b, right? The a is the slope, and the b the y-intercept. So, if you have to couples \[(x_{1}, y_{1}) and (x_{2},y_{2})\], you get : \[a=\frac{ y_{2}-y_{1} }{ x_{2}-x_{1} }\] Afterward, you want to find the value of b, which is pretty easy. You choose one of your couple, and you plug you values into your equation y=ax+b. In your case, you'll get, example: -3=a(1)+b (your a has a value that you find earlier). So, you only have to isolate b in your equation and you win :)

OpenStudy (anonymous):

So is my y value 4.5? or 3/2 in slope form? Now that I have that, how would I find the x?

OpenStudy (texaschic101):

first we find the slope using the slope formula : slope(m) = (y2 - y1) / (x2 - x1) (-5,3) x1 = -5 and y1 = 3 (1,-3) x2 = 1 and y2 = -3 now we sub slope(m) = (-3 - 3) / (1 - (-5) slope(m) = -6/6 slope(m) = -1 now we will use y = mx + b form and find b because we already know the other terms you can use either of your set of points... (-5,3) x = -5 and y = 3 slope(m) = -1 now we sub y = mx + b 3 = -1(-5) + b 3 = 5 + b 3 - 5 = b -2 = b so our equation is : y = -x - 2 To find the x intercept, sub in 0 for y y = -x - 2 0 = -x - 2 x = -2 your x coordinate of the x intercept (-2,0) is -2

OpenStudy (anonymous):

Wow, it was a bad simple math mistake on my part when getting the slope, after that, thanks for the info and help! I got it now!

OpenStudy (texaschic101):

good to hear :)